If a boat is traveling downstream to a destination and it takes 40 minutes in a 3 mph current. Traveling the same distance going upstream with 3mph current it takes 60 minutes. How fast would the boat be traveling in still water?


I'm not even sure how to go about this but I think going downhill the boat is going 13.3mph and uphill the boat is having to go 20mph?

Answer by mananth(16949)

Boat speed =x mph
current speed 3 mph

time with current 0.67 hours
time against current 1 hours
0.67(x+3)=1(x-3)
0.67x+ 2.00 =1x-3
1 x -0.67 x = 2 + 3
0.33 x = 5
x = 15 mph Boat speed